OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 3(x + 8)2(x - 8)3 -8, multiplicity 2; 8, multiplicity 3 4, multiplicity 1; 8, multiplicity 1; -8, multiplicity 1 -8, multiplicity 3; 8, multiplicity 2 4, multiplicity 1; -8, multiplicity 3; 8, multiplicity 3 I think it's C.

OpenStudy (anonymous):

**f(x)=3(x+8)^2(x-8)^3

OpenStudy (e.mccormick):

First, is that \(f(x) = 3(x + 8)^2(x - 8)^3\)

OpenStudy (anonymous):

yeah

OpenStudy (e.mccormick):

Then what you do to solve these is set each factor to 0 and solve for x to find the 0s, then the power is the multiplicity. Watc your signs when doing this.

OpenStudy (anonymous):

wait what do i set to zero? the numbers or the x's?

OpenStudy (e.mccormick):

The factors with a variable.

OpenStudy (e.mccormick):

\(f(x)=(x-a)^m(x+b)^n\) fins 0s and multiplicity example: For \((x-a)^m\) \(x-a=0\) \(x-a+a=0+a\) \(x=a\) with a multiplicity of \(m\). For \((x+b)^n\) \(x+b=0\) \(x+b-b=0-b\) \(x=-b\) with a multiplicity of \(n\).

OpenStudy (anonymous):

oh i got it

OpenStudy (e.mccormick):

So do you see what is wrong with C?